a purse contains $6.35 in quarters and nickles. The number of quarters is one less than 2 times the number of nickles. Find the number of each coin in the purse.
Answer by hamsanash1981@gmail.com(151)

Let the number of nickels be x, then quarters is 2x-1.
Total Nickels = .05x and total quarters is .25(2x-1) = .5x -.25
.05x+0.5x-0.25 = 6.35
.55x = 6.35+.25
.55x = 6.6
x = 6.6/0.55
x = 660/55
x = 60/5
x = 12
quarters = 23 and nickels = 12